# Few-shot Object Localization

### v1

@@ III-A3 Gradient Branch
where $p_{k}\in\{(-1,0),(0,-1),(0,0),(0,1),(1,0)\}$ when $N=9$, since only the gradient information in the horizontal and vertical directions is aggregated. $\theta\in[0,1]$ is used to balance the intensity-level and gradient-level information.
In our framework, both the query and support features undergo processing by a single CCD-C(HV) layer with shared parameters. Specifically, the query feature $F_{Q}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and support feature $F_{S}\in\mathbb{R}^{1\times 256\times 3\times 3}$ are optimized to enhanced features $F_{Q}^{C}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and $F_{S}^{C}\in\mathbb{R}^{1\times 256\times 3\times 3}$ after being processed by CCD-C layer individually. This approach not only reduces computational cost but also boosts the model’s sensitivity to gradient information, such as edges and brightness. In essence, it enables the model to detect boundaries and changes in brightness more effectively across different samples, thereby enhancing accuracy in object identification and categorization. This serves to alleviate the issue of object omission triggered by gradient-related factors such as brightness and edges.
@@
After processing by the aforementioned two branches, both the support and query features are enriched with gradient and deformation information. This enhancement equips our model to efficiently and accurately decipher the rich yet noisy information encapsulated within these features. However, enhancing them separately overlooks the holistic nature of the features and fails to leverage the inherent relationships between the information they contain. To tackle these issues, we leverage 3D convolution to integrate information from both branches and facilitate comparison between the support and query features. 3D convolution allows for simultaneous processing of information across multiple channels by introducing a depth dimension. Thus, for a given position $p$ on the input feature map $x$, the corresponding value at the same position $p$ on the output feature map $y$ can be mathematically expressed as:
@@
where $D$ represents the depth dimension of the input feature map, $p_{k}\in\{(-1,-1,-1),(-1,-1,0),\ldots,(0,1,1),(1,1,1)\\ \}$. We concatenate the outputs of the two branches along different dimensions, effectively creating a multi-dimensional feature space. This concatenated feature is then fed into a 3D convolutional operation. Here, the support features are sequentially stacked and used as the convolution kernel, convolving with the similarly stacked query features. Throughout this process, the convolution along the depth dimension integrates both deformation and gradient information, thereby generating an association map between the support and query features. This approach results in a more comprehensive and accurate feature representation.
Specifically, the model inputs the DC-enhanced query feature $F_{Q}^{D}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and support feature $F_{S}^{D}\in\mathbb{R}^{1\times 256\times 3\times 3}$, as well as CCD-C-enhanced features $F_{Q}^{C}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and $F_{S}^{C}\in\mathbb{R}^{1\times 256\times 3\times 3}$. Subsequently, the 3D convolution employs the parameters of the support features $F_{S}^{D}$ and $F_{S}^{C}$ as convolution kernel weights. It then conducts convolution in the $H$, $W$, and depth $D$ directions, corresponding to the stacking direction of $F_{Q}^{D}$ and $F_{Q}^{C}$, finally obtaining the preliminary similarity map $S\in\mathbb{R}^{1\times 256\times 128\times 128}$. Since the stacking order in the depth direction is specified during input, the support features outputted by the two branches convolve on the query feature without interfering with each other. This process can be represented by the following formula:
@@
III-C Self-query Attention
The DFA module and 3D convolution significantly enhance the model’s feature extraction capability. However, they solely focus on mining the similarity between support samples and query images to generate a similarity map, disregarding the distribution information of objects in the image. To overcome this limitation, inspired by the self-support matching [17], we integrate the distribution information of the query image into the obtained similarity map. This integrated information guides the optimization process, rather than relying solely on a blind search for similar samples in the similarity map. Essentially, the model gradually learns to utilize the information from the query image to steer optimization of the obtained feature comparison results. As a result, it achieves high precision in feature matching even with limited information from support samples.
Figure 5: Demonstration of Self-query (SQ) Module. The SQ module receives input from the similarity map $S$, which is generated by the Dual-path Feature Augmentation module, and the original query map feature $F_{Q}$. Initially, it employs a shared convolution layer to introduce non-linearity to both $S$ and $F_{Q}$. Subsequently, it computes the cosine similarity between $S$ and $F_{Q}$, resulting in the self-query weights $W$ that connect them. It then performs element-wise addition between $W$ and $S$, which means to guide the distribution information of $S$ through the $F_{Q}$. Finally, after passing through another convolution layer, the SQ module produces the optimized similarity map denoted as $S_{SQ}$
To this end, the Self-query (SQ) module firstly uses two 1$\times$1 convolution layers with shared parameters to introduce non-linearity to the original query features and the obtained similarity map. It then uses cosine similarity comparison to obtain the similarity weight, then perform element-wise addition between this weight and the similarity map for optimization. In detail, the query feature hints the distribution of all foreground objects and the background on the query image, while the similarity map describes the similarity of the query image relative to the support sample, that is, the difference between all query image pixels and the support sample. We use cosine similarity to compare the activated query features and the similarity map, essentially leveraging the distribution information of the objects in the query image to guide the similarity map, guiding the model to pay more attention to the interesting parts in the query features and the similarity map. Then, the model uses this similarity as weights to optimize the similarity map, making the foreground objects sharper and the background smoother, so that the similarity map can more significantly reflect the support samples on the query image. The overall process is shown in Fig. 5.
Aforementioned process can be represented as:
@@ III-C Self-query Attention
where $In\_Conv$ and $Out\_Conv$ represents 1$\times$1 convolution, $S\in\mathbb{R}^{1\times 256\times 128\times 128}$ and $F_{Q}\in\mathbb{R}^{1\times 256\times 128\times 128}$ represents the similarity map outputted by 3D convolution and original feature extracted from query image respectively. They are both inputted to a shared convolution layer and calculated a similarity weight $W\in\mathbb{R}^{1\times 128\times 128}$ by cosine similarity. Since the shape of the similarity map $S$ and similarity weight $W$ is same, thus they can directly perform element-wise addition and obtain the SQ-enhanced similarity map $S_{SQ}^{T}\in\mathbb{R}^{1\times 256\times 128\times 128}$. Then, $S_{SQ}^{T}$ is processed by another 1$\times$1 convolution layer to get final enhanced similarity map $S_{SQ}\in\mathbb{R}^{1\times 256\times 128\times 128}$ as the output of our proposed SQ module.
To be more specific, the aforementioned process can be represented with the following pseudo algorithm:
Algorithm 1 Self-query Algorithm
@@
1:  function $Self\_query(F_{Q},S)$
@@
6:  Self-query similarity map $S_{SQ}$
Here, $F_{Q}^{T}$, $S^{T}$ and $S_{SQ}^{T}$ represents the intermediate variables of $F_{Q}$, $S$ and $S_{SQ}$, respectively. The preliminary similarity map $S$ obtained by 3D convolution, after prompted by the distribution information of query image, can more significantly reflect the interesting samples in the query image with its own information. Therefore, it can detect more tricky objects and maintain competitive localization capabilities even when the support information is scarce.
III-D Post Processing
After optimizing the similarity map, the model post-processes it to yield the final localization results. Our post-processing, which is divided into two steps, begins with regressing the similarity map. We first construct a regression head using a sequence of convolution layers, followed by Leaky ReLU activation and bi-linear upsampling layers using designs of [35, 37, 40]. Then, we use the Focal Inverse Distance Transform (FIDT) map as the regression target [18]. In our FSOL task, distinguishing dense and overlapping objects is more challenging than in ordinary object localization tasks due to the small number of samples. The FIDT map is more dispersed compared to other types of density maps, making it more suitable for dense localization tasks, and therefore more suitable for FSOL tasks.
After optimizing the localization density map for dense objects, following [18], we use the Local-Maxima-Detection-Strategy (LMDS) method to predict the local maximum values of the density map. In FIDTM, given a series of $M$ candidate points, only points with values greater than threshold $T_{a}=100/255\cdot M$ are selected as positive samples to filter out false positives. In FSOL tasks, object sizes vary greatly across different datasets, so we divide the datasets into dense datasets and sparse datasets based on the average sample sizes. For dense datasets, we set the threshold to $T_{a}=40/255\cdot M$. For sparse datasets, because the object size is often larger than the objects in dense datasets, we modify this threshold to $T_{a}=60/255\cdot M$. At the same time, we set a lower limit of 0.06 for all candidate point values to filter out noise, meaning that all candidate points with values less than 0.06 are considered negative samples.
@@ IV-A Datasets
In this paper, we mainly run experiments on four datasets.
@@ IV-A1 FSC-147
The FSC-147 dataset is a unique collection of images designed for the task of few-shot counting [37]. It encompasses a wide variety of 147 different object categories, ranging from everyday items like kitchen utensils and office stationery, to more complex entities such as vehicles and animals. This dataset contains 6135 images while the number of objects in each image vary significantly, with a range from as few as 7 to as many as 3731 objects. Each object instance in an image is marked with a dot at its approximate center. Moreover, for each image, three object instances are randomly chosen as exemplar instances and are annotated with axis-aligned bounding boxes.
@@ IV-A2 ShangHaiTech
The ShangHaiTech dataset is a comprehensive crowd counting dataset that consists of two parts: ShangHaiTech Part A and ShangHaiTech Part B [21]. ShangHaiTech Part A is a collection of 482 images that were gathered from the internet. These images are diverse and cover a wide range of scenarios, making them ideal for training robust crowd counting models. The dataset is further divided into a training subset with 300 images and a testing subset with 182 images. On the other hand, ShangHaiTech Part B is a set of 716 images that were captured on the bustling streets of Shanghai. This part of the dataset provides a realistic representation of real-world crowd scenarios. It is also split into a training subset with 400 images and a testing subset with 316 images. Each image in both parts of the dataset is annotated with a dot at the center of each person’s head, providing precise ground truth labels for crowd counting. In total, the dataset contains annotations for 330,165 individuals.
@@ IV-A3 CARPK
The car parking lot dataset (CARPK) is for object counting and localization tasks, specifically designed for counting cars in parking lots [41]. The dataset contains 1448 images of nearly 90,000 cars, each annotated with a bounding box. The training set contains 3 scenes, while another scene is used for test.
@@ IV-B Metrics
IV-B1 Localization Metrics
This paper uses the euclidean distance between the center point of the predicted object and the ground truth point to represent the localization error, which can be represented as:
@@ IV-B1 Localization Metrics
To better compare the localization performance of different models, a threshold $\sigma$ is set, which represents the maximum allowable error. For each predicted point, we calculate its localization error with all ground truth points. If the minimum error is less than or equal to $\sigma$, then we consider this predicted point a positive sample, indicating correct object location by the model. Otherwise, this predicted point is considered a negative sample, indicating that the model has incorrectly located the object. FSOL is a very challenging task, with a large span of sample sizes. Therefore, this paper mainly uses $\sigma_{l}=10$ as the high threshold when studying model performance and uses $\sigma_{l}=8$ on specific datasets to align with other model results. Specifically, $\sigma_{l}=10$ means that if the distance between a predicted point and Ground truth point is less than or equal to 10 pixels, then this predicted point is a positive sample. At the same time, this paper presents results using 4 and 5 as the low threshold $\sigma_{s}$ to demonstrate the model’s localization performance under stricter conditions.
The main metrics used in this paper include precision, recall, and F1-score. Precision measures how many of the points predicted by the model as positive samples are true positive samples; recall measures how many of all true positive samples are successfully detected by the model. Precision and recall can be represented by the following formulas:
@@
where $TP$ represents the model correctly matching the predicted points and ground truth points; $FN$ represents the model not matching any predicted points for ground truth points; $FP$ represents the points predicted by the model do not have corresponding ground truth points.
The F1-score is a measure of the model’s accuracy on binary classification systems, which classify examples as ‘positive’ or ‘negative’. The F1 score is the harmonic mean of precision and recall, with a range from 0 to 1. The closer it is to 1, the better the model’s classification performance. The F1 score can be represented by the following formula:
@@
IV-B2 Counting Metrics
We additionally use Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) to evaluate the model’s counting ability, indicating that our model has excellent generalization on other tasks. The definitions of MAE and RMSE are as follows:
@@ IV-B2 Counting Metrics
where $N$ is the number of test images, $y_{i}$ is the number of ground truth points in the i-th image, and $\hat{y}_{i}$ is the number of predicted points in the i-th image. The closer MAE and RMSE are to 0, the closer the model’s prediction results are to the ground truth results.

### v3

@@ III-A3 Gradient Branch
where $p_{k}\in\{(-1,0),(0,-1),(0,0),(0,1),(1,0)\}$, since only the gradient information in the horizontal and vertical directions is aggregated. $\theta\in[0,1]$ is used to balance the intensity-level and gradient-level information.
In our framework, both the query and support features are processed by a single CCD-C(HV) layer with shared parameters. Specifically, the query feature $F_{Q}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and support feature $F_{S}\in\mathbb{R}^{1\times 256\times 3\times 3}$ are optimized to enhanced features $F_{Q}^{C}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and $F_{S}^{C}\in\mathbb{R}^{1\times 256\times 3\times 3}$ after being processed by CCD-C layer individually. This approach not only reduces computational cost but also boosts the model’s sensitivity to gradient information, such as edges and brightness. In essence, it enables the model to detect boundaries and brightness changes more effectively across different samples, thereby enhancing accuracy in object identification and categorization. This serves to alleviate the issue of object omission triggered by gradient-related factors such as brightness and edges.
@@
After processed by the aforementioned two branches, both the support and query features are enriched with gradient and deformation information. This enhancement equips our model to efficiently and accurately learn the rich yet noisy information encapsulated within these features. However, enhancing them separately overlooks the holistic nature of the features and fails to leverage the inherent relationships between the information they contain. To tackle these issues, we leverage 3D convolution to integrate information from both branches and facilitate comparison between the support and query features. 3D convolution allows for simultaneous processing of information across multiple channels by introducing a depth dimension. Thus, for a given position $p$ on the input feature map $x$, the corresponding value at the same position $p$ on the output feature map $y$ can be mathematically expressed as:
@@
where $D$ represents the depth dimension of the input feature map, $p_{k}\in\{(-1,-1,-1),(-1,-1,0),\ldots,(0,1,1),(1,1,1)\\ \}$. We concatenate the outputs of the two branches along different dimensions, effectively creating a multi-dimensional feature space. This concatenated feature is then subjected to a 3D convolutional operation. Here, the support features are sequentially stacked and used as the convolution kernel, convolving with the similarly stacked query features. Throughout this process, the convolution along the depth dimension integrates both deformation and gradient information, thereby generating an association map between the support and query features. In other words, enhanced features in different aspects can be aggregated together to yield a more comprehensive and accurate feature representation.
Specifically, the model inputs the DC-enhanced query feature $F_{Q}^{D}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and support feature $F_{S}^{D}\in\mathbb{R}^{1\times 256\times 3\times 3}$, as well as CCD-C-enhanced features $F_{Q}^{C}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and $F_{S}^{C}\in\mathbb{R}^{1\times 256\times 3\times 3}$. Subsequently, the 3D convolution employs the parameters of the support features $F_{S}^{D}$ and $F_{S}^{C}$ as convolution kernel weights. It then conducts convolution in the $H$, $W$, and depth $D$ directions, corresponding to the stacking direction of $F_{Q}^{D}$ and $F_{Q}^{C}$, finally obtaining the preliminary similarity map $S\in\mathbb{R}^{1\times 256\times 128\times 128}$. Since the stacking order in the depth direction is specified during input, the support features outputted by the two branches convolve on the corresponding query feature without interfering with each other. This process can be represented by the following formula:
@@
III-C Self Query Module
The DFA module and 3D convolution significantly enhance the model’s feature extraction capability. However, they solely focus on mining the similarity between support samples and query images to generate a similarity map, disregarding the distribution information of objects in the query image. To overcome this limitation, inspired by the self-support matching [17], we design the Self Query (SQ) module to integrate the distribution information of the query image into the obtained similarity map. This integrated information guides the optimization process, rather than relying solely on a blind search for similar samples in the similarity map. Essentially, the model gradually learns to utilize the information from the query image to steer optimization of the obtained feature comparison results. As a result, it achieves high precision in feature matching even with limited information from support samples.
Figure 5: Demonstration of Self Query (SQ) module. The SQ module enhances the model’s perception of the object distribution by integrating information from the similarity map $S$ and the original query image features $F_{Q}$. Initially, it applies a shared convolution layer to both $S$ and $F_{Q}$, thereby introducing non-linearity and capturing similar patterns. Next, it calculates the cosine similarity between $S$ and $F_{Q}$ to obtain self query weights $W$. These weights are added element-wise to $S$, enabling the distribution information from $S$ to guide optimization through $F_{Q}$. Finally, after passing through another convolution layer, the SQ module generates the optimized similarity map $S_{SQ}$.
The SQ module enhances model’s perception of the object distribution by integrating query image features and the similarity map through several key steps. Initially, it employs two 1$\times$1 convolution layers with shared parameters to introduce non-linearity and capture similar patterns in both the original query features and the obtained similarity map. Following the computation of cosine similarity between the query features and the similarity map, self query weights are derived. These weights are then added element-wise to the original query features, allowing the model to leverage the distribution information of objects in the query image for optimizing the similarity map. This iterative process enables the model to prioritize relevant regions within both the query features and the similarity map. Subsequently, the resulting similarity is employed as weights to refine the similarity map, enhancing the clarity of foreground objects while maintaining smoothness in the background. This refined representation enables the similarity map to more accurately depict support samples present on the query image. The comprehensive process is illustrated in Fig. 5. The aforementioned process can be represented as:
@@ III-C Self Query Module
where $In\_Conv$ and $Out\_Conv$ represents 1$\times$1 convolution, $S\in\mathbb{R}^{1\times 256\times 128\times 128}$ and $F_{Q}\in\mathbb{R}^{1\times 256\times 128\times 128}$ represents the similarity map outputted by 3D convolution and original feature extracted from query image respectively. They are both inputted to a shared convolution layer and calculated a similarity weight $W\in\mathbb{R}^{1\times 128\times 128}$ by cosine similarity. Since the shape of the similarity map $S$ and similarity weight $W$ is same, thus they can directly perform element-wise addition and obtain the SQ-enhanced similarity map $S_{SQ}^{T}\in\mathbb{R}^{1\times 256\times 128\times 128}$. Then, $S_{SQ}^{T}$ is processed by another 1$\times$1 convolution layer to get final enhanced similarity map $S_{SQ}\in\mathbb{R}^{1\times 256\times 128\times 128}$ as the output of our SQ module.
Algorithm 1 Self Query Algorithm
@@
1:  function $Self\ \ Query(F_{Q},S)$
@@
6:  Self Query similarity map $S_{SQ}$
To be more specific, we present the process using Algorithm 1. Here, $F_{Q}^{T}$, $S^{T}$ and $S_{SQ}^{T}$ denote the intermediate variables of $F_{Q}$, $S$ and $S_{SQ}$, respectively. The initial similarity map $S$, obtained through 3D convolution, is sensitized by the distribution information from the query image. This adaptation enables $S$ to more accurately represent the distinctive features present in the query image. Consequently, it becomes adept at identifying complex objects, maintaining robust localization capabilities even in scenarios where support information is limited.
III-D Regression and Post Processing
After obtaining the optimized similarity map, the model undergoes regression and post-processing steps. The regression process involves upsampling the feature map to match the size of the input image, facilitating subsequent loss computation or further post-processing to derive coordinate information. This regression step is achieved through a series of convolutional layers, followed by Leaky ReLU activation and bilinear upsampling layers, drawing inspiration from designs presented in [33, 35, 38]. Consequently, a location map of identical dimensions to the original image is generated. During training, the location map is directly compared with ground truth to compute the loss, subsequently updating the model parameters. During testing, the location map undergoes post-processing to extract target position information.
Specifically, the post-processing is implemented based on the Local-Maxima-Detection-Strategy described in FIDTM  [18]. Given a series of $M$ candidate points, only points with values greater than the threshold $T_{a}=100/255$ are selected as positive samples to filter out false positives. In FSOL tasks, object sizes vary significantly across different datasets. Therefore, we categorize datasets into dense and sparse datasets based on the average sample sizes. For dense datasets, the threshold is set to $T_{a}=40/255$. However, for sparse datasets, where object sizes tend to be larger compared to dense datasets, we adjust this threshold to $T_{a}=60/255$. Additionally, we enforce a lower limit of 0.06 for all candidate point values to filter out noise, implying that candidate points with values less than 0.06 are considered negative samples.
@@ IV-A Datasets
To validate the effectiveness of our method and demonstrate the high performance of the baseline framework, this paper primarily conducts experiments on four datasets. These datasets include dedicated few-shot datasets as well as densely populated datasets of crowds and vehicles.
@@ IV-A1 FSC-147
The FSC-147 dataset is a unique collection of images tailored for the task of few-shot counting [35]. It features a diverse array of 147 distinct object categories, spanning from everyday items like kitchen utensils and office stationery to more complex entities such as vehicles and animals. This dataset comprises 6135 images, with the number of objects in each image varying significantly, ranging from as few as 7 to as many as 3731 objects. Each object instance within an image is denoted by a dot at its approximate center. Furthermore, for each image, three object instances are randomly selected as exemplar instances and annotated with axis-aligned bounding boxes.
@@ IV-A2 ShangHaiTech
The ShangHaiTech dataset is a comprehensive crowd counting dataset comprising two parts: ShangHaiTech Part A and ShangHaiTech Part B [21]. ShangHaiTech Part A consists of 482 images sourced from the internet. These images exhibit diversity and encompass a wide range of scenarios, rendering them suitable for training robust crowd counting models. The dataset is further partitioned into a training subset comprising 300 images and a testing subset comprising 182 images.
Conversely, ShangHaiTech Part B comprises 716 images captured on the bustling streets of Shanghai. This segment of the dataset offers a realistic depiction of real-world crowd scenarios. It is divided into a training subset with 400 images and a testing subset with 316 images. Each image in both parts of the dataset is annotated with a dot marking the center of each person’s head, providing precise ground truth labels for crowd counting. In total, the dataset contains annotations for 330,165 individuals.
@@ IV-A3 CARPK
The car parking lot dataset (CARPK) is tailored for object counting and localization tasks, focusing specifically on counting cars in parking lots [39]. This dataset comprises 1448 images featuring nearly 90,000 cars, with each car annotated by a bounding box. The training set consists of three scenes, while an additional scene is reserved for testing purposes.
@@ IV-B Metrics
IV-B1 Localization
This paper quantifies the localization error using the Euclidean distance between the center point of the predicted object and the corresponding ground truth point. Mathematically, it can be expressed as:
@@ IV-B1 Localization
To facilitate a comprehensive comparison of localization performance among different models, a threshold $\sigma$ is established, representing the maximum allowable error. For each predicted point, its localization error with all ground truth points is computed. If the minimum error is less than or equal to $\sigma$, the predicted point is deemed a positive sample, indicating correct object localization by the model. Otherwise, if the minimum error exceeds $\sigma$, the predicted point is classified as a negative sample, signifying an incorrect object localization by the model. Given the wide range of sample sizes in FSOL, this paper predominantly utilizes $\sigma_{l}=10$ as the high threshold to assess model performance. However, on specific datasets, $\sigma_{l}=8$ is adopted to align with other model results. Furthermore, this paper presents results using 4 and 5 as the low threshold $\sigma_{s}$ to demonstrate the model’s localization performance under stricter conditions.
The primary localization metrics employed in this paper include precision, recall, and F1-score. Precision assesses the proportion of predicted positive samples that are true positive samples, while recall evaluates the proportion of all true positive samples that are successfully detected by the model. These metrics can be represented by the following formulas:
@@
where $TP$ denotes true positives, indicating instances where the model correctly matches predicted points with ground truth points; $FN$ represents false negatives, indicating cases where the model fails to match any predicted points for ground truth points; $FP$ signifies false positives, indicating instances where the points predicted by the model lack corresponding ground truth points.
The F1-score, a measure of model accuracy in binary classification, combines precision and recall into a single metric. It ranges from 0 to 1, with higher scores indicating better performance. The F1-score is calculated using the harmonic mean of precision and recall. Mathematically, it can be expressed as:
@@
IV-B2 Counting
In addition, we use Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) to evaluate the model’s counting ability, indicating that our model has excellent generalization on other tasks. The definitions of MAE and RMSE are as follows:
@@ IV-B2 Counting
where $N$ represents the total number of test images, $y_{i}$ denotes the number of ground truth points in the i-th image, and $\hat{y}_{i}$ represents the number of predicted points in the i-th image. Lower values of MAE and RMSE indicate better alignment between the model’s predictions and the ground truth results.